# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v3

@@ A.3 Hypernetwork
Hypernetworks, which are neural networks designed to generate the weights of other networks, allow for dynamic adjustment of model parameters. This adaptability enables the model to better suit different tasks and datasets, thereby improving overall performance. By utilizing hypernetworks, the number of models that need to be trained individually can be significantly reduced. Traditional methods require separate models for each task, whereas hypernetworks can generate weights for multiple tasks. This capability enhances training efficiency. In our task, we use hypernetworks to generate adapters for the question and input, which are then inserted into the model. This helps the model incorporate the knowledge targeted by the question, corresponding to implicit imagination. Compared to traditional efficient fine-tuning, this process is more aligned with the goal of imagination.
@@ B.1 Background
Our task formulation follows retrieval augmented models for QA Guu et al. (2020); Sachan et al. (2021). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(q,a,c)$ comprising question $q$, answer $a$, and context $c$, where $q,a,c\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function, represented as $p(a|q)$, wherein the models decode a string $a$ that serves as an abstractive answer to a given query $q$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{a}=\arg\max_{a\in\mathcal{V}^{*}}p(a|q,\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to omitting a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(a|q,c)$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(a|q,c)$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(a|q,c)=\sum\limits^{c_{i}\in c}\limits_{c\in TopK(q)}p(a|q,c_{i})p(c_{i}|q)$, where $Topk(q)$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $q$.
@@ B.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit imagination of the context generator to imagine a short dummy useful document is:
@@
IMcQA-l 10 49.92 60.03 25.79 34.35 69.67 20.28 30.19 54.94 51.52
Table 6: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance. IMcQA attains optimal performance both in-distribution and OOD under diverse document configurations.
@@ B.3 Implementations
We consider conducting experiments using three different sizes of T5, namely T5-base, T5-large, T5-3b, and Llama2-7B, Llama2-13B Touvron et al. (2023). Due to memory constraints and online distillation limitations, A100 supports processing 20 documents for T5-3b, while Llama2 does not support distillation. All experiments with T5-3b are conducted on 2 A100 GPUs, T5-large on 2 A6000 GPUs, and T5-Base on 2 RTX 3090 GPUs. However, experiments with Llama2-7b and 13b, except for IMcQA on 2 A100 GPUs, are tested on 8 RTX 3090 GPUs.
@@ C.1 Supervised Performance
Additionally, in comparison with LoRA Hu et al. (2021) methods, IMcQA enhances EM scores by an average of +2.2%. In the closed-book setting, the LoRA method manifests a substantial decrease in performance, likely attributable to the inadequacy of learning sufficient knowledge via questions for storage in the LoRA module. On the other hand, IMcQA harnesses both explicit and implicit imaginations to exploit knowledge for improved outcomes. These results indicate that the knowledge stored in the LLMs’ parameters can still be further exploited.
@@ C.2 OOD Results
Table 6 shows the full OOD results in QA. It can be observed that our method has the best OOD generalization ability on all three benchmarks. Although LoRA performs well on the in-distribution part, its performance is generally poor on OOD, with some even showing negative performance. This highlights the importance of the domain adaptability of the implicit imagination Hypernetwork in our method, which generates LoRA adapter weights based on input.
@@ C.3 Zero-Shot Results
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with increased document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text imagination during the compression phase.
@@ C.4 OOD and Ablation Experiment Results
Here, we supplement the experimental results of LoRA and IMcQA under supervised fine-tuning in closed-book settings and the ablation results of feedforward neural network (FFN) and Long Context Distillation (LCD). It can be observed that our method like LoRA, belongs to parameter-efficient fine-tuning, and because we share the Hypernetwork to generate LoRA adapter weights, we fine-tune fewer parameters.
From Table 9, it can be seen that releasing FFN can bring more performance improvement, possibly because adding LoRA in Attention cannot fully utilize enough knowledge Yao et al. (2022). With the support of LCD, performance is further improved, with an average increase in EM of +5%. This also proves the effectiveness of our proposed LCD. In comparison with IMcQA and LoRA, it becomes more evident that LoRA tends to transfer knowledge to the LoRA module, resulting in low generalization. Our method enhances knowledge activation through dynamic generation, showing significant effects not only in-distribution but also in OOD.
@@ C.5 Error Analysis
Table 9: OOD and ablation experiment results. * denotes the results are from the existing papers and LCD denotes Long Context Distillation.
@@
Imagine Please write a high-quality answer for the given question using your knowledge and the provided imagined compressed results (some of which might be irrelevant). Only give me the answer and do not output any other words. Imagined Context: {context} Answer the question based on your knowledge and the given imagined context. Question: {question} Answer:
@@
IMcQA (Ours) 220M 0 23.89 22.69 30.31
IMcQA-l (Ours) 770M 0 29.32 30.11 32.68
IMcQA-xl (Ours) 3b 0 29.59 35.71 37.40
@@
IMcQA (Ours) 220M 10 47.01 64.95 46.36
IMcQA-l (Ours) 770M 10 49.92 69.67 51.52
IMcQA-xl (Ours) 3b 5 50.87 70.34 52.78
IMcQA-l (Ours) 770M 30 53.1 70.5 52.0
@@
IMcQA (Ours) 220M 10 46.22 66.70 51.43
IMcQA-l (Ours) 770M 10 48.83 70.85 54.52
IMcQA-xl (Ours) 3b 5 49.23 72.18 55.39
# Imagination Augmented Generation (Ours)

### v4

@@ A.3 Hypernetwork
Hypernetworks Ha et al. (2016) offer a solution that reduces the dependency on gradient descent for specific domains. Methods such as Hypertuning Phang et al. (2022) and HINT Ivison et al. (2023) use hypernetworks to transform inputs into parameter-efficient modules, thereby reducing computation and enhancing model generalization.
Hypernetworks, which are neural networks designed to generate the weights of other networks, allow for dynamic adjustment of model parameters. This adaptability enables the model to better suit different tasks and datasets, thereby improving overall performance. By utilizing hypernetworks, the number of models that need to be trained individually can be significantly reduced. Traditional methods require separate models for each task, whereas hypernetworks can generate weights for multiple tasks. This capability enhances training efficiency. In our task, we use hypernetworks to generate adapters for the question and input, which are then inserted into the model. This helps the model incorporate the knowledge targeted by the question, corresponding to implicit awakening. Compared to traditional efficient fine-tuning, this process is more aligned with the goal of awakening.
@@ B.1 Background
Our task formulation follows retrieval augmented models for QA Guu et al. (2020); Sachan et al. (2021). Let $\mathcal{V}^{*}$ denote the infinite set, encompassing all potential strings over the tokens in vocabulary $\mathcal{V}$, and this includes the empty string. An instance within a QA dataset is defined as a triplet $(\boldsymbol{q},\boldsymbol{a},\boldsymbol{c})$ comprising question $\boldsymbol{q}$, answer $\boldsymbol{a}$, and context $c$, where $\boldsymbol{q},\boldsymbol{a},\boldsymbol{c}\in\mathcal{V}^{*}$. Conventionally, the context $c$ is drawn from the knowledge corpus $\mathcal{Z}$, like Wikipedia, whereby $\mathcal{Z}\subset\mathcal{V}^{*}$.
The goal of QA is to learn a distribution function, represented as $p(\boldsymbol{a}|\boldsymbol{q})$, wherein the models decode a string $\boldsymbol{a}$ that serves as an abstractive answer to a given query $\boldsymbol{q}$. In a closed-book setting, LLMs directly encode the given question and predict the answer Roberts et al. (2020b). Specifically, considering the context $c$ as the empty string, the reliance is solely on the model parameters, i.e., $\hat{\boldsymbol{a}}=\arg\max_{\boldsymbol{a}\in\mathcal{V}^{*}}p(\boldsymbol{a}|\boldsymbol{q},\theta)$, where $\theta$ represents the LLMs’ parameters. However, employing a direct approach of requesting models to output answers frequently results in subpar performance, primarily attributable to omitting a substantial amount of world knowledge during the process. Therefore, a popular approach is open domain setting, which marginalizes $p(\boldsymbol{a}|\boldsymbol{q},\boldsymbol{c})$ over contexts $c$ in the knowledge corpus Lewis et al. (2020); Sachan et al. (2021) or generated from models Yu et al. (2023). Given the computational infeasibility of calculating probabilities for all contexts, $p(\boldsymbol{a}|\boldsymbol{q},\boldsymbol{c})$ is approximated to the sum of probabilities for top $k$ contexts, i.e., $p(\boldsymbol{a}|\boldsymbol{q},\boldsymbol{c})=\sum\limits^{\boldsymbol{c}_{i}\in\boldsymbol{c}}\limits_{c\in\text{Topk}(\boldsymbol{q})}p(\boldsymbol{a}|\boldsymbol{q},\boldsymbol{c}_{i})p(\boldsymbol{c}_{i}|\boldsymbol{q})$, where $\text{Topk}(\boldsymbol{q})$ denotes the set of resulting top $k$ passages after the retrieval or generated with a query $\boldsymbol{q}$.
@@ B.2 Prompts for Explicit Imagine with LLMs
The prompt for explicit awakening of the context generator to imagine a short dummy useful document is:
@@
AAG-l 10 49.92 60.03 25.79 34.35 69.67 20.28 30.19 54.94 51.52
Table 7: OOD results. The primary row in the table header delineates the dataset trained, while the underscored secondary row demonstrates the in-distribution performance. AAG attains optimal performance both in-distribution and OOD under diverse document configurations.
@@ B.3 Implementations
We consider conducting experiments using three different sizes of T5, namely T5-base, T5-large, T5-3b, and Llama2-7B, Llama2-13B Touvron et al. (2023). Due to memory constraints and online distillation limitations, A100 supports processing 20 documents for T5-3b, while Llama2 does not support distillation. All experiments with T5-3b are conducted on 2 A100 GPUs, T5-large on 2 A6000 GPUs, and T5-Base on 2 RTX 3090 GPUs. However, experiments with Llama2-7b and 13b, except for AAG on 2 A100 GPUs, are tested on 8 RTX 3090 GPUs.
@@ C.1 Supervised Performance
Additionally, in comparison with LoRA Hu et al. (2021) methods, AAG enhances EM scores by an average of +2.2%. In the closed-book setting, the LoRA method manifests a substantial decrease in performance, likely attributable to the inadequacy of learning sufficient knowledge via questions for storage in the LoRA module. On the other hand, AAG harnesses both explicit and implicit awakenings to exploit knowledge for improved outcomes. These results indicate that the knowledge stored in the LLMs’ parameters can still be further exploited.
@@ C.2 OOD Results
Table 7 shows the full OOD results in QA. It can be observed that our method has the best OOD generalization ability on all three benchmarks. Although LoRA performs well on the in-distribution part, its performance is generally poor on OOD, with some even showing negative performance. This highlights the importance of the domain adaptability of the implicit awakening Hypernetwork in our method, which generates LoRA adapter weights based on input.
@@ C.3 Zero-Shot Results
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with increased document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text awakening during the compression phase.
@@ C.4 OOD and Ablation Experiment Results
Here, we supplement the experimental results of LoRA and AAG under supervised fine-tuning in closed-book settings and the ablation results of feedforward neural network (FFN) and Long Context Distillation (LCD). It can be observed that our method like LoRA, belongs to parameter-efficient fine-tuning, and because we share the Hypernetwork to generate LoRA adapter weights, we fine-tune fewer parameters.
From Table 12, it can be seen that releasing FFN can bring more performance improvement, possibly because adding LoRA in Attention cannot fully utilize enough knowledge Yao et al. (2022). With the support of LCD, performance is further improved, with an average increase in EM of +5%. This also proves the effectiveness of our proposed LCD. In comparison with AAG and LoRA, it becomes more evident that LoRA tends to transfer knowledge to the LoRA module, resulting in low generalization. Our method enhances knowledge activation through dynamic generation, showing significant effects not only ID but also in OOD.
@@ C.5 Error Analysis
C.6 Number of Document Compression
In the first stage, we sampled 30,000 instances from the training sets of NQ and TQA, respectively, and used all 3,417 instances from the WebQ training set. To determine the number of retrieved documents to use for each dataset in stage 1, we conducted tests using the FiD (T5-Base) experiment. As shown in Table 10, we can find that compressing five documents yielded relatively good performance. Consequently, we decided to compress five documents for each instance.
1 5 10 20 50
FiD 34.69 41.27 - - 46.59
# Document Compression
AAG 32.57 38.19 35.17 32.12 36.83
Table 10: Performance Metrics for Different Configurations
Appendix D Case Study
This study illustrates the differences in how three paradigms—RAG, GAG, and AAG—utilize documents during reasoning as shown in Table 11. RAG retrieves ten documents from an external knowledge base, while GAG employs ChatGPT to generate ten documents with higher similarity. For illustration, we present only the content of the first document. Conversely, AAG uses its proprietary context generator to create virtual compressed documents containing more information. The token counts in parentheses show that AAG requires significantly fewer tokens for processing compared to the other two methods, thus enhancing inference speed and reducing computational overhead.
Question: what style of art did henri matisse use?
Method # Docs Documents
RAG 10 Henri Matisse Henri Émile Benoît Matisse (; 31 December 1869 – 3 November 1954) was a French artist, known for both his use of colour and his fluid and original draughtsmanship. He was a draughtsman, printmaker, and sculptor, but is known primarily as a painter. Matisse is commonly regarded, along with Pablo Picasso, as one of the artists who best helped to define the revolutionary developments in the visual arts throughout the opening decades of the twentieth century, responsible for significant developments in painting and sculpture. The intense colorism of the works he painted between 1900 and 1905 brought him… (1860 tokens)
GAG 10 Henri Matisse is considered one of the most important artists of the 20th century. He is known for his use of color and his distinctive style of painting and sculpture. Matisse was a member of the French avant-garde movement and his work was influenced by other artists such as Paul C0̆0e9zanne and Pablo Picasso… (1540 tokens)
AAG(ours) 1 Henri Matisse context: a period of convalescence attack of appendic. He discovered ä kind of paradise as he later described it, to become an deeply disappointing his. In 19 he returned to study art the Acad0̆0e9mie and became a student of William-Adolphe Bouguereau Gustave Moreau Initially he painted still lif and landscapes a traditional style at which reasonable proficiency Mat was influenced the works earlier masters such as Jean-Bapt-Sim0̆0e9on Ch Nicolas Pous Watteau, as well artists, such as 0̆0c9douard Manet a body of work spanning over a half-century, won him recognition as a leading figure in modern art. Matisse was born in Le Cateau-Cambr0̆0e9sis, in the Nord department in northern France, the oldest son of a prosperous grain merchant. He grew up in Bohain-en-Vermandois, Picardie, France. In 1887 he went to Paris to study law, working as a court administrator in Le Cateau-Cambr0̆0e9sis after gaining his qualification. He first started to paint in 1889, after his mother brought him art supplies (280 tokens)
Table 11: A inference case with used documents from WQ.
@@ Appendix D Case Study
Table 12: OOD and ablation experiment results in closed-book setting. * denotes the results are from the existing papers and LCD denotes Long Context Distillation.
@@
Awakening Please write a high-quality answer for the given question using your knowledge and the provided imagined compressed results (some of which might be irrelevant). Only give me the answer and do not output any other words. Generated Context: {context} Answer the question based on your knowledge and the given generated context. Question: {question} Answer:
@@
AAG (Ours) 220M 0 23.89 22.69 30.31
AAG-l (Ours) 770M 0 29.32 30.11 32.68
AAG-xl (Ours) 3b 0 29.59 35.71 37.40
@@
AAG (Ours) 220M 10 47.01 64.95 46.36
AAG-l (Ours) 770M 10 49.92 69.67 51.52
AAG-xl (Ours) 3b 5 50.87 70.34 52.78
AAG-l (Ours) 770M 30 53.1 70.5 52.0
@@
AAG (Ours) 220M 10 46.22 66.70 51.43
AAG-l (Ours) 770M 10 48.83 70.85 54.52
AAG-xl (Ours) 3b 5 49.23 72.18 55.39
# Awakening Augmented Generation (Ours)